# On the survival of the long-lived inner disk of PDS 70 Key Words.: accretion, accretion disks – circumstellar matter – planets and satellites: formation – protoplanetary disks1 Introduction

### 11 Mar 2024 v1

@@ 3.2 Comparison with ALMA observations
Figure 4 shows the comparison of models and current ALMA observations of PDS 70. For this comparison, synthetic images are created at the same wavelength (855 m) as the observations presented in Benisty et al. (2021), assuming the dust density distribution from the models shown in Fig. 1 after 1 Myr of evolution. The beam for the synthetic images is assumed to be as in the ALMA observations, i.e., 0.046”0.36” with a position angle of 51.5.
@@
The synthetic images in Fig 4 demonstrate that the dust evolution models assuming cm or cm do not reproduce the emission of the inner disk that is observed. This is also seen in the left panel of Fig 5, which corresponds to the comparison of the azimuthally averaged radial intensity profiles of the deprojected images from the observations and the models.
@@
The model in which cm can well reproduce the emission of the inner disk, even if the total dust mass within this region is very low (). It is worth mentioning that the models with small grains also predict an inner shoulder just within the main ring. This emission is located near the location of the outer planet, in which a small bump is created in between the two gaps. There is a similar shape of emission in the actual observations (Band 7; Benisty et al. 2021) with an inner ‘shoulder’ within the bright ring (Fig. 4 and Fig 5), however, the models underpredict the flux of such structure because the grain size of the dust particles is small at that location. Nonetheless, it is interesting to see that the emission of such ‘shoulder’ also increases as the minimum grain size in the simulations is smaller and when the filtration is less efficient (left panel in Fig 5). Higher angular resolution are required to understand what are the properties of such emission, and test if for example it is a well separated ring or not.
@@
The right panel of Fig 5 shows the averaged radial intensity profiles of deprojected images obtained at different ALMA bands, showing how the emission decreases at longer wavelengths in the model with cm and assuming the same angular resolution than the observations from Benisty et al. (2021) (left panel in Fig. 4). The detection of the inner disk would depend on the sensitivity of the observations to detect it. Already in Band 3 (3 mm) the flux of the inner disk is of few micron-Jy per beam.
@@
Figure 6 shows the total millimetre flux as a function of wavelength from the same ALMA bands of the right panel of Fig. 4. The fluxes are calculated at three different regions (as denoted in the right panel of Fig. 5), the inner disk (16 au), the gap ( 36 au), and the ring (112 au). Using curvefit within scipy (Virtanen et al. 2020), we fit to find the spectral index of each region. The spectral index of the ring has values of . Because the total flux is dominated by the ring emission, the spectral index integrated over the entire disk is also 3.2, which is similar to the value within the gap ().
@@
In this model, the spectral index is slightly higher in the inner disk, with a value of . Hence, these models suggest that the spectral index that can be obtained from observations using optically thin wavelengths will show a high spectral index of with a slighly increased in the inner disk. This value of the spectral index is high compared to the averaged values found in protoplanetary disks (e.g., Tazzari et al. 2021), however, for a set of transition disks Pinilla et al. (2014) showed that the observed spectral index increased with the size of the cavity, and that overall transition disks have higher spectral index than the disks without cavities. Specifically, the linear relation between the spectral index and the cavity size (defined as the peak observed in millimeter emission) is . Using a cavity size of au (Keppler et al. 2019), the expected is 3.2 as found in our models suggesting that the spatially integrated spectral index of PDS 70 agrees with observations of other transition disks.
@@
These values, however, do not match the spectral index of 2.7 obtained from the total flux at 855 m (176 mJy, Keppler et al. 2019) and at 1.3,mm (57 mJy, Facchini et al. 2021), which could be due to the high optical depth of the observations, specially at 855 m. Observations at multiple wavelengths where the emission is optically thin are needed to test current models and the results of Fig. 6.
@@ 4.1 Effect of the gas disk mass and turbulence
The coupling of dust particles to the gas is quantified by the Stokes number given in Eq. 3. Particles with St1 are well coupled to the gas and can be dragged along with it, while St1 are well trapped and hence filtered out. There is an uncertainty related to the bulk density of the dust particles, and hence an uncertainty related to the dust composition and porosity. In this work, we tested dust evolution models, in which the bulk density was assumed to be 1.6 g cm, similar to the value obtained when the DSHARP opacities are assumed. This change leads to Stokes numbers that are a factor of 2 higher for the same grain size and gas surface density, which does not affect significantly the results presented in this work. In our models, the main uncertainty in the Stokes number is the gas surface density that directly depends on the gas disk mass, which still remains unknown from observations of most disks and, PDS 70 is not the exception.
@@
Portilla-Revelo et al. (2023) performed detailed thermochemical models of the ALMA Band 6 observations, in particular to model the distribution of three CO isotopologues, in order to constrain the gas distribution of PDS 70. Their analysis suggests a gas mass of within the dust depleted gap (between 16 and 41 au). The gas mass within the same region in our models is in good agreement with the values from Portilla-Revelo et al. (2023).
@@
The total disk mass (inside 130 au) from Portilla-Revelo et al. (2023) is , while in our models it is . Assuming a lower disk mass would imply that the Stokes number is higher for a given grain size, and as result less filtration is expected. In other words, a less massive disk would imply that the minimum grain size would need to be lower than the best case scenario explored in this work (i.e., cm). Such nano-sized particles have been suggested to be present from observations of a couple of protoplanetary disks (Habart et al. 2021; Kokoulina et al. 2021).
@@
It is still possible that the gas disk mass obtained from CO observations is undestimated because of different chemical and physical processes that can convert most CO into other refractory molecules or because CO is sequestrated into icy bodies (Schwarz et al. 2018; Krijt et al. 2020). As a test, we performed three dust evolution models where the gas disk mass is increased by a factor of ten and assuming the same initial grain size than in the models presented in Sect. 3. The summary of these simulations are presented in Fig. 7, which corresponds to the evolution of the dust mass for each case (as Fig. 2). These results show how when the disk mass is higher, the filtration is lower for a given initial particle size distribution. For instance, for the models with cm, the dust mass from 1 to 10 Myr is (similar to the dust mass in the models of Sect. 3, but for cm) and does not continuously decrease as in the case of Fig. 2. This highlights the importance of obtaining better constraints of the gas disk mass, which can have a large impact on the dust filtration effect and the dust density distribution in the inner disk of PDS 70.
@@
Recently, Trapman et al. (2022) suggested a new method to measure gas disk masses from observing NH and CO isotopologues, which has been tested against HD observations of a couple of disks. Currently, these observations have not been done for PDS 70, but it would be interesting to test in future observations such a method and compare it with the disk masses from Portilla-Revelo et al. (2023), that are in agreement with the values used in this work.
@@
The level of turbulence can also change the effect of filtration, with higher turbulence implying less filtration. However, multiple models and observations of disks suggest that turbulence is rather low (, see e.g. the recent review by Rosotti 2023). For transition disks in particular, a higher viscosity would imply a less empty or no cavity at all independent of the mass of the planets that are embedded in the disk (de Juan Ovelar et al. 2016). On the other hand, a lower viscosity for PDS 70 disk would imply more filtration, but also multiple rings and gaps created by the planets that are embedded in this disk (e.g., Dong et al. 2017; Bae et al. 2018). As a result, a viscosity of , as used in this work seems appropiate for PDS 70. The intrinsic nature of viscosity is in anycase complex, it depends on how the disks are ionized and coupled to the magnetic fields (e.g., Lesur et al. 2023), which can lead to spatial variations of depending on different enviromental, stellar, and disk properties (Delage et al. 2021).
@@ 4.2 Inner emission and comparison with JWST observations
JWST/MIRI observations of PDS 70 revealed the presence of water in its inner disk (au, Perotti et al. 2023). Specifically, the HO spectrum is best fitted with a slab model of gas at 600 K inside 0.05 au, and a column density of cm. Small grains are of great importance to explain this water emission because of two reasons: first, small grains can help to shield ultraviolet light and prevent HO photodissociation (Heays et al. 2017), and second because they can be the particles that can travel across the gap and enrich the inner disk, as demonstrated in this work.
We calculate the total dust mass that is potentially reaching the snowline in PDS 70 by taking the dust loss rates () at the inner edge of the simulations in Sect. 2, and integrating over either 0-1 Myr, where the vast majority of the accreted dust comes from the initial material that was inside the gap location, or over 1-3.5 Myr where most of the material comes from the potential replenishment from the outer to the inner disk. We do not integrate over the entire 10 Myr of the dust evolution models because the age of PDS 70 is approximately 5.4 Myr, and as discussed before the initial time in the dust evolution models should be seen as at least 2 Myr of the disk age.
@@
where is the mass of a water molecule, i.e. 18 g/mol, and is the disk inclination taken to be 51.7. Assuming that is 10% or 90%, we obtain that is cm when in 10% and cm when in 90%. There are several assumptions in this calculation, in particular we are neglecting any chemical effect that can change the water abundance, but it is just to illustrate that potentially the water column density observed with JWST could solely originate from drifting particles that come from the outer disk, which need to be small enough in our models (0.1 m) to pass through the gap formed by the two giant planets in this system.
@@ 5 Conclusions
PDS 70 remains as the best laboratory to study the influence of the formation of giant planets on the disk structure and the potential formation of terrestrial planets within the system. One of the intriguing properties of PDS 70 is the existence of an inner disk, revealed by the spatially resolved emission observed with ALMA at different wavelengths. Its survival is puzzling because dust evolution models suggest that complete filtering of dust particles by the gap created by the planets would stop the dust replenishment from the outer to the inner disk. We confirm this result by performing state-of-the art dust evolution models, when the minumum grain size in the models is larger than 0.1m. As a consequence, the initial dust that was within the planets’ orbit eventually migrates inwards and the inner disk is depleted within the first million-year of dust evolution.
To maintain an inner disk, the minimum grain size in the models need to be lower than 0.1m. These results mainly depend on the assumption taken for the gas disk mass, and we performed models with values that are in agreement with recent thermochemical models that are compared with ALMA observations (Portilla-Revelo et al. 2023). Only when grains are that small, they are diffused and dragged along with the gas throughout the planets’ gap. These small grains grow and drift in the inner disk, but the constant reservoir of dust particles that are trapped in the outer edge of the gap and that are continuously fragmenting helps to preserve the inner disk on million-year timescales. Our flux predictions at millimetre wavelength of these models agree with current ALMA observations. Other possible uncertainties originate from the planet masses. If the planet masses are lower than Jupiter mass, the filtration effect is not as efficient as presented in this work. However, observations and models of PDS 70 disk robustly support that PDS 70b and PDS 70c are more massive than Jupiter.
It is possible that the water emission observed recently with JWST from the inner disk of PDS 70 originates from these small grains that cross the gap and then drift towards the inner disk, reaching the snowline. This idea can be a mirror of our Solar System where Jupiter and Saturn formation could have blocked the majority of pebbles in the outer disk, but where still water ice could have been carried by the very small grains in the parental disk of the Solar System. Future observations that give better constraints on the gas disk mass and the level of turbulence in PDS 70 will help to test current models.
@@ Appendix A Radial resolution test
We perform a dust evolution model to test the radial resolution of our simulations. The left panel of Fig. 8 shows the dust density distribution at 1 Myr of evolution for the case of cm as in the top-left panel of Fig. 1 but with a higher number of radial cells in the simulations (). The right panel of Fig. 8 shows the total dust disk mass as a function of time for both values of the radial resolution ( and , due to the computational cost of the high resolution simulations, we only show results up to 1.5 Myr after the dust disk mass does not change anymore over time) Figure 8 show no differences in the results when assuming vs.

### 21 Mar 2024 v2

@@ 3.2 Comparison with ALMA observations
Figure 4 shows the comparison of models and current ALMA observations of PDS 70. For this comparison, synthetic images were created at the same wavelength (855 m) as the observations presented in Benisty et al. (2021), assuming the dust density distribution from the models shown in Fig. 1 after 1 Myr of evolution. The beam for the synthetic images was assumed to be as in the ALMA observations, that is, 0.046”0.36”, with a position angle of 51.5.
@@
The synthetic images in Fig 4 demonstrate that the dust evolution models assuming cm or cm do not reproduce the emission of the inner disk that is observed. This is also shown in the left panel of Fig 5, which corresponds to the comparison of the azimuthally averaged radial intensity profiles of the deprojected images from the observations and the models.
@@
The model in which cm can reproduce the emission of the inner disk well even when the total dust mass within this region is very low (). It is worth mentioning that the models with small grains also predict an inner shoulder just within the main ring. This emission is located near the location of the outer planet, in which a small bump is created in between the two gaps. There is a similar shape of emission in the actual observations (Band 7; Benisty et al. 2021) with an inner shoulder within the bright ring (Fig. 4 and Fig 5), but the models underpredict the flux of this structure because the grain size of the dust particles is small at that location. Nonetheless, it is interesting to see that the emission of this shoulder also increases as the minimum grain size in the simulations is smaller and when the filtration is less efficient (left panel in Fig 5). A higher angular resolution is required to understand the properties of this emission and test whether it is a well-separated ring, for example.
@@
The right panel of Fig 5 shows the averaged radial intensity profiles of deprojected images obtained at different ALMA bands, showing that the emission decreases at longer wavelengths in the model with cm and assuming the same angular resolution as for the observations from Benisty et al. (2021) (left panel in Fig. 4). The detection of the inner disk would depend on the sensitivity of the observations to detect it. Already in Band 3 (3 mm), the flux of the inner disk is a few micron-Jy per beam.
@@
Figure 6 shows the total millimeter flux as a function of wavelength from the same ALMA bands as in the right panel of Fig. 4. The fluxes were calculated in three different regions (as denoted in the right panel of Fig. 5), the inner disk (16 au), the gap ( 36 au), and the ring (112 au). Using curvefit within scipy (Virtanen et al. 2020), we fit to find the spectral index of each region. The spectral index of the ring has values of . Because the total flux is dominated by the ring emission, the spectral index integrated over the entire disk is also 3.2, which is similar to the value within the gap ().
@@
In this model, the spectral index is slightly higher in the inner disk, with a value of . Hence, these models suggest that the spectral index that can be obtained from observations using optically thin wavelengths will show a high spectral index of with a slightly increased inner disk. This value of the spectral index is high compared to the averaged values found in protoplanetary disks (e.g., Tazzari et al. 2021), but for a set of transition disks, Pinilla et al. (2014) showed that the observed spectral index increased with the size of the cavity and that overall, transition disks have a higher spectral index than the disks without cavities. Specifically, the linear relation between the spectral index and the cavity size (defined as the peak observed in millimeter emission) is . For a cavity size of au (Keppler et al. 2019), the expected is 3.2, as found in our models, suggesting that the spatially integrated spectral index of PDS 70 agrees with observations of other transition disks.
@@
These values, however, do not match the spectral index of 2.7 obtained from the total flux at 855 m (176 mJy, Keppler et al. 2019) and at 1.3,mm (57 mJy, Facchini et al. 2021), which could be due to the high optical depth of the observations, especially at 855 m. Observations at multiple wavelengths where the emission is optically thin are needed to test current models and the results of Fig. 6.
@@ 4.1 Effect of the gas disk mass and turbulence
The coupling of dust particles to the gas is quantified by the Stokes number given in Eq. 3. Particles with St1 are well coupled to the gas and can be dragged along with it, while St1 are well trapped and hence filtered out. There is an uncertainty related to the bulk density of the dust particles, and hence an uncertainty related to the dust composition and porosity. In this work, we tested dust evolution models in which the bulk density was assumed to be 1.6 g cm, similar to the value obtained when the DSHARP opacities are assumed. This change leads to Stokes numbers that are higher by a factor of 2 for the same grain size and gas surface density, which does not affect the results presented in this work significantly. In our models, the main uncertainty in the Stokes number is the gas surface density, which directly depends on the gas disk mass. This still remains unknown from observations of most disks and, PDS 70 is no exception.
@@
Portilla-Revelo et al. (2023) performed detailed thermochemical models of the ALMA Band 6 observations, in particular, to model the distribution of three CO isotopologs, in order to constrain the gas distribution of PDS 70. Their analysis suggested a gas mass of within the dust-depleted gap (between 16 and 41 au). The gas mass within the same region in our models is , which agrees well with the values from Portilla-Revelo et al. (2023).
@@
The total disk mass (inside 130 au) from Portilla-Revelo et al. (2023) is , while in our models, it is . An assumed lower disk mass would imply that the Stokes number is higher for a given grain size, and as result, less filtration is expected. In other words, a less massive disk would imply that the minimum grain size would need to be lower than the best-case scenario explored in this work (i.e., cm). Nano-size particles like this have been suggested to be present from observations of some protoplanetary disks (Habart et al. 2021; Kokoulina et al. 2021).
@@
It is still possible that the gas disk mass obtained from CO observations is undestimated because of different chemical and physical processes that can convert most CO into other refractory molecules or because CO is sequestered into icy bodies (Schwarz et al. 2018; Krijt et al. 2020). As a test, we performed three dust evolution models in which the gas disk mass was increased by a factor of ten and assumed the same initial grain size as in the models presented in Sect. 3. The summary of these simulations is presented in Fig. 7, and it corresponds to the evolution of the dust mass for each case (as Fig. 2). These results show that when the disk mass is higher, the filtration is lower for a given initial particle size distribution. For instance, for the models with cm, the dust mass from 1 to 10 Myr is (similar to the dust mass in the models of Sect. 3, but for cm) and does not continuously decrease as in the case of Fig. 2. This highlights the importance of obtaining better constraints on the gas disk mass, which can have a large impact on the dust filtration effect and the dust density distribution in the inner disk of PDS 70.
@@
Recently, Trapman et al. (2022) suggested a new method for measuring gas disk masses from observing NH and CO isotopologs, which was tested against HD observations of several disks. These observations have not been performed for PDS 70so far, but it would be interesting to test this method in future observations and compare it with the disk masses from Portilla-Revelo et al. (2023), which agree with the values used in this work.
@@
The level of turbulence can also change the effect of filtration, with higher turbulence implying less filtration. However, multiple models and observations of disks suggest that turbulence is rather low (, see e.g. the recent review by Rosotti 2023). For transition disks in particular, a higher viscosity would imply a less empty cavity or no cavity at all, regardless of the mass of the planets that are embedded in the disk (de Juan Ovelar et al. 2016). On the other hand, a lower viscosity for PDS 70 disk would imply more filtration, but also multiple rings and gaps created by the planets that are embedded in this disk (e.g., Dong et al. 2017; Bae et al. 2018). As a result, a viscosity of , as we used here, seems appropriate for PDS 70. The intrinsic nature of viscosity is complex in any case. It depends on how the disks are ionized and coupled to the magnetic fields (e.g., Lesur et al. 2023), which can lead to spatial variations of depending on different enviromental, stellar, and disk properties (Delage et al. 2021).
@@ 4.2 Inner emission and comparison with JWST observations
The JWST/MIRI observations of PDS 70 revealed the presence of water in its inner disk (au, Perotti et al. 2023). Specifically, the HO spectrum is best fit with a slab model of gas at 600 K inside 0.05 au and a column density of cm. Small grains are of great importance to explain this water emission for two reasons: First, small grains can help to shield ultraviolet light and prevent HO photodissociation (Heays et al. 2017), and second, they can be the particles that can travel across the gap and enrich the inner disk, as demonstrated in this work.
We calculated the total dust mass that potentially reaches the snowline in PDS 70 by taking the dust-loss rates () at the inner edge of the simulations in Sect. 2, and integrating over either 0-1 Myr, where the vast majority of the accreted dust comes from the initial material that was inside the gap location, or over 1-3.5 Myr, where most of the material comes from the potential replenishment from the outer to the inner disk. We did not integrate over the entire 10 Myr of the dust evolution models because the age of PDS 70 is approximately 5.4 Myr, and as discussed before, the initial time in the dust evolution models should be seen as at least 2 Myr of the disk age.
@@
where is the mass of a water molecule, that is, 18 g/mol, and is the disk inclination taken to be 51.7. Assuming that is 10% or 90%, we obtain that is cm when is 10% and cm when is 90%. There are several assumptions in this calculation. In particular, we neglected any chemical effect that can change the water abundance, but this was done just to illustrate that the water column density observed with JWST might solely originate from drifting particles that come from the outer disk, which need to be small enough in our models (0.1 m) to pass through the gap formed by the two giant planets in this system.
@@ 5 Conclusions
PDS 70 remains the best laboratory for studying the influence of the formation of giant planets on the disk structure and the potential formation of terrestrial planets within the system. One of the intriguing properties of PDS 70 is its inner disk, which was revealed by the spatially resolved emission observed with ALMA at different wavelengths. Its survival is puzzling because dust evolution models suggest that complete filtering of dust particles by the gap created by the planets would stop the dust replenishment from the outer to the inner disk. We confirmed this result by performing advanced dust evolution models, in which the minumum grain size in the models was larger than 0.1m. As a consequence, the initial dust that was within the orbit of the planets eventually migrated inward, and the inner disk was depleted within the first million years of the dust evolution.
To maintain an inner disk, the minimum grain size in the models needs to be lower than 0.1m. These results mainly depend on the assumption taken for the gas disk mass, and we performed models with values that agreed with recent thermochemical models, which were compared with ALMA observations (Portilla-Revelo et al. 2023). Only when the grains are that small are they diffused and dragged along with the gas throughout the gap of the planets. These small grains grow and drift in the inner disk, but the constant reservoir of dust particles that are trapped in the outer edge of the gap and that fragment continuously helps to preserve the inner disk on million-year timescales. Our flux predictions at millimeter wavelength of these models agree with current ALMA observations. Other possible uncertainties originate from the planet masses. If the planet masses are lower than the mass of Jupiter, the filtration effect is not as efficient as presented in this work. However, observations and models of the PDS 70 disk robustly support that PDS 70b and PDS 70c are more massive than Jupiter.
It is possible that the water emission that was recently observed with the JWST from the inner disk of PDS 70 originates from these small grains that cross the gap and then drift toward the inner disk, where they reach the snowline. This idea can be a mirror of our Solar System, in which the formation of Jupiter and Saturn could have blocked the majority of pebbles in the outer disk, but where water ice could still have been carried by the very small grains in the parental disk of the Solar System. Future observations that better constrain the gas disk mass and the level of turbulence in PDS 70 will help us to test our models.
@@ Appendix A Radial resolution test
We performed a dust evolution model to test the radial resolution of our simulations. The left panel of Fig. 8 shows the dust density distribution at 1 Myr of evolution for the case of cm as in the top left panel of Fig. 1, but with a higher number of radial cells in the simulations (). The right panel of Fig. 8 shows the total dust disk mass as a function of time for both values of the radial resolution ( and ; due to the computational cost of the high-resolution simulations, we only show results up to 1.5 Myr after the dust disk mass no longer changes over time). Figure 8 shows no differences in the results when assuming versus